# SWAT: Sliding Window Adversarial Training for Gradual Domain Adaptation

### v1

@@ 5.5 Ablation Study
Ablation Study on Adversarial Flow Matching
We validate the necessity of hierarchical feature learning through controlled experiments on Rotated MNIST. By progressively enabling multi-scale feature aggregation in our sliding window framework, we observe systematic performance improvements across 2–4 domain configurations (Fig. LABEL:fig:abl-conv). The shallowest configuration (equivalent to raw pixel statistics without Adversarial train) underperforms by over 25% compared to our Adversarial Flow Matching, confirming the insufficiency of low-level features.
Ablation Study on Adversarial Training
This study evaluates the impact of sliding window adversarial training (SWAT) mechanism across three datasets: Rotated MNIST, Color-Shift MNIST, and Portraits. The key comparison is between configurations that do not use adversarial methods (NO), and the sliding window adversarial training (SWAT).
@@
In the Rotated MNIST dataset (Fig. LABEL:fig:abl-subfig1), the accuracy without any adversarial alignment (NO) drops significantly with inter-domain steps, whereas incorporating SWAT with adversarial flow matching improves accuracy. For the Color-Shift MNIST dataset (Fig. LABEL:fig:abl-subfig2), SWAT significantly enhances accuracy, achieving near-optimal performance across inter-domain steps. In the Portraits dataset (Fig. LABEL:fig:abl-subfig3), SWAT outperforms the baseline NO method and any previous static transport methods.
6 Conclustion
This study advances domain adaptation through the novel Sliding Window Adversarial Training (SWAT) framework, which addresses critical limitations in handling gradual domain shifts. By integrating adversarial learning with flow matching, SWAT eliminates dependency on predefined intermediate domains and enables continuous feature flow alignment via bidirectional adversarial optimization. Extensive validation across diverse benchmarks demonstrates SWAT’s superiority over existing methods, with significant improvements in domain-invariant feature preservation and adaptation stability. This work establishes a foundation for exploring more sophisticated domain adaptation strategies in complex real-world scenarios, paving the way for enhanced generalization capabilities in evolving environments.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Implementation
Our transport architecture includes generators composed of a single residual block containing three linear layers. The discriminator is built with three linear layers, each having 128 hidden units and paired with ReLU activation functions. We used the Adam optimizer for optimization (Kingma & Ba, 2014), Dropout for regularization (Srivastava et al., 2014), and Batch Normalization to stabilize training (Ioffe & Szegedy, 2015). The number of intermediate domains generated between source and target domains is treated as a hyperparameter, with the model’s performance evaluated for 0, 1, 2, 3, or 4 intermediate domains. All the code was ran on NVIDIA RTX 4090 GPUs.
In addition, we followed (Kumar et al., 2020) to filter out the 10% of data points where the model’s predictions exhibit the least confidence. However, instead of relying on the typical uncertainty measure, we define the confidence level as the difference between the largest and the second-largest values in the model’s output. We have found that this produces better results and we use this setting in all comparative tests.
@@
D.1D.2D.3D.4D.5D.620202020404040406060606080808080100100100100Domain (D.)Accuracy (%)Accuracy vs Domain Shift for Different DatasetsRotated MNISTPortraitsColor-Shift MNISTCover Type
@@ A.2 Results of Our Method
We present a comparison of our proposed SWAT method with multiple datasets, including Rotated MNIST, Color-Shift MNIST, Portraits and Cover Type, as detailed in Tables 4 through 7. Each experiment was repeated multiple times, with the results shown as mean values along with variance intervals. The leftmost column of each table represents the performance obtained using only adversarial training, which corresponds to the method without flow matching.
In Tables 4 to 7, the column ”# Given Domains” indicates the number of domains included in the experiment, comprising both the source and the target domains. The ”Inter-domain counts in SWAT” columns indicate the number of inter-domain steps taken between the given domains in the dataset. The entire process is equivalent to including (”# Given Domains - 1”) × (”# Inter-domain counts in SWAT + 1”) + 1 training step, which includes self-training of GAN and the encoder $f$ and classifier $g$. For example, with four domains and three intermediate steps, the total number of training steps is calculated as (4 - 1) × (3 + 1) + 1 = 13 small steps.
@@
3 $97.9\pm 1.9$ $99.4\pm 0.1$ $99.4\pm 0.0$ $99.2\pm 0.4$ $\mathbf{99.5\pm 0.0}$
4 $99.4\pm 0.0$ $\mathbf{99.6\pm 0.0}$ $99.5\pm 0.0$ $99.5\pm 0.1$ $\mathbf{99.6\pm 0.0}$
5 $99.5\pm 0.0$ $\mathbf{99.6\pm 0.0}$ $99.5\pm 0.1$ $99.4\pm 0.3$ $99.5\pm 0.1$
6 $\mathbf{99.6\pm 0.0}$ $99.4\pm 0.3$ $99.2\pm 0.5$ $99.4\pm 0.1$ $99.5\pm 0.1$
@@
It is important to highlight that the highest accuracy points are typically found in the upper-right and lower-left corners of the table. This suggests that as the number of given domains increases, the adversarial flow matching tends to become more complete, eliminating the need for additional intermediate steps to refine the flow. This observation demonstrates that our method of constructing flows matching between domains is particularly effective when only a few domains are given, and the sliding window adversarial training is highly effective all the time.

### v2

@@ 5.4 Ablation Study
Continuous Feature Flow
We conduct ablation study of continuous feature flow on Rotated MNIST. By progressively enabling multi-scale feature aggregation in our sliding window framework, we observe significant performance improvements across 2–4 domain settings (Fig. 6(a)). The shallowest setting, corresponding to a shallow neural network without adversarial training, performs over 25% worse than our feature flow matching approach, highlighting the limitations of low-level features in capturing transferable representations.
(a) Feature Extraction Strategies
(b) Rotated MNIST
(c) Color-Shift MNIST
(d) Portraits
Figure 6: Ablation analysis of SWAT. (a) Comparison of flow matching with feature extraction strategies. (b-d) Performance of the SWAT method on Rotated MNIST, Color-Shift MNIST, and Portrait datasets (with 3 intermediate domains), showing the accuracy changes of different training strategies (NO: no adversarial, GST[7], GOAT[13] with 0-4 Inter-domain step counts.
Bidirectional Flow Matching
@@
In the Rotated MNIST dataset (Fig. 6(b)), the accuracy without any adversarial alignment (NO) drops significantly with inter-domain steps, whereas incorporating SWAT with feature flow matching improves accuracy. For the Color-Shift MNIST dataset (Fig. 6(c)), SWAT significantly enhances accuracy, achieving near-optimal performance across inter-domain steps. In the Portraits dataset (Fig. 6(d)), SWAT outperforms the baseline NO method and any previous static transport methods.
Table 4: Analysis of different change strategies of pp. "Ours" denotes gradually increasing $p$ at equal intervals, "Fixed" keeps $p$ a constant value of 0.5, “Rand” samples $p$ randomly from a uniform distribution $U(0,1)$ at each step, and “Sorted” uses a fixed set of random values in ascending order.
Methods
Rotated MNIST Portraits
0 1 2 3 4 0 1 2 3 4
Ours $83.3\text{\scriptsize{$\pm$0.9}}$ $85.0\text{\scriptsize{$\pm$0.5}}$ $\mathbf{86.1\text{\scriptsize{$\pm$0.4}}}$ $\mathbf{86.9\text{\scriptsize{$\pm$0.2}}}$ $\mathbf{88.1\text{\scriptsize{$\pm$1.5}}}$ $82.9\text{\scriptsize{$\pm$1.2}}$ $\mathbf{84.6\text{\scriptsize{$\pm$0.2}}}$ $\mathbf{85.0\text{\scriptsize{$\pm$0.9}}}$ $\mathbf{85.1\text{\scriptsize{$\pm$0.2}}}$ $\mathbf{85.3\text{\scriptsize{$\pm$0.1}}}$
Sorted $\mathbf{84.1\text{\scriptsize{$\pm$0.8}}}$ $\mathbf{86.4\text{\scriptsize{$\pm$0.6}}}$ $86.0\text{\scriptsize{$\pm$1.7}}$ $86.3\text{\scriptsize{$\pm$0.1}}$ $85.7\text{\scriptsize{$\pm$0.5}}$ $82.7\text{\scriptsize{$\pm$0.5}}$ $84.0\text{\scriptsize{$\pm$0.6}}$ $84.2\text{\scriptsize{$\pm$0.1}}$ $84.3\text{\scriptsize{$\pm$0.2}}$ $84.5\text{\scriptsize{$\pm$0.1}}$
Rand $83.4\text{\scriptsize{$\pm$0.2}}$ $80.9\text{\scriptsize{$\pm$7.0}}$ $84.5\text{\scriptsize{$\pm$2.8}}$ $86.3\text{\scriptsize{$\pm$0.9}}$ $86.1\text{\scriptsize{$\pm$0.4}}$ $82.4\text{\scriptsize{$\pm$0.5}}$ $84.0\text{\scriptsize{$\pm$0.6}}$ $84.2\text{\scriptsize{$\pm$0.2}}$ $84.3\text{\scriptsize{$\pm$0.2}}$ $84.1\text{\scriptsize{$\pm$0.1}}$
Fixed $83.3\text{\scriptsize{$\pm$0.0}}$ $83.7\text{\scriptsize{$\pm$0.0}}$ $83.8\text{\scriptsize{$\pm$0.1}}$ $83.8\text{\scriptsize{$\pm$0.1}}$ $84.1\text{\scriptsize{$\pm$0.0}}$ $\mathbf{83.9\text{\scriptsize{$\pm$0.3}}}$ $83.4\text{\scriptsize{$\pm$0.0}}$ $83.1\text{\scriptsize{$\pm$2.1}}$ $84.1\text{\scriptsize{$\pm$0.1}}$ $84.7\text{\scriptsize{$\pm$0.3}}$
Sliding Window Mechanism
Table 4 reports results across five inter-domain adaptation steps. The sliding window mechanism (“Ours”) consistently achieves the best average accuracy, e.g., 88.1% vs. 84.1% (Fixed) and 86.1% (Rand) at step 4 on Rotated MNIST with substantially lower standard deviation (±0.5 vs. ±7.0 at step 1). These results confirm that gradually increasing $p$ produces more stable and optimal adaptation than holding $p$ fixed, sampling it at random, or reordering random draws.
6 Conclusion
This work proposes a sliding window mechanism to improve the adversarial training, which splits large domain shifts into multiple micro transfers through local, dynamic and continuous feature alignment, enabling fine-grained distribution matching. Building upon this training paradigm, we present the Sliding Window Adversarial Training (SWAT), a novel framework for GDA that incorporates the sliding window mechanism with adversarial flow matching to enable continuous and stable feature alignment. Extensive experimental results demonstrate the superior effectiveness and robustness of SWAT across diverse benchmarks.
@@ A.1 Implementation
Our transport architecture includes generators composed of a single residual block containing three linear layers. The discriminator is built with three linear layers, each having 128 hidden units and paired with ReLU activation functions. We used the Adam optimizer for optimization [50], Dropout for regularization [51], and Batch Normalization to stabilize training [52]. The number of intermediate domains generated between source and target domains is treated as a hyperparameter, with the model’s performance evaluated for 0, 1, 2, 3, or 4 intermediate domains. All the code was ran on NVIDIA RTX 4090 GPUs.
In addition, we followed [7] to filter out the 10% of data points where the model’s predictions exhibit the least confidence. However, instead of relying on the typical uncertainty measure, we define the confidence level as the difference between the largest and the second-largest values in the model’s output. We have found that this produces better results and we use this setting in all comparative tests.
@@
D.1D.2D.3D.4D.5D.62020404060608080100100Domain (D.)Accuracy (%)Accuracy vs Domain Shift for Different DatasetsRotated MNISTPortraitsColor-Shift MNISTCover Type
@@ A.2 Results of Our Method
We present a comparison of our proposed SWAT method with multiple datasets, including Rotated MNIST, Color-Shift MNIST, Portraits, and Cover Type, as detailed in Tables 5 through 8. Each experiment was repeated multiple times, with the results shown as mean values along with variance intervals. The leftmost column of each table represents the performance obtained using only adversarial training, which corresponds to the method without flow matching.
In Tables 5 to 8, the column "# Given Domains" indicates the number of domains included in the experiment, comprising both the source and the target domains. The "Inter-domain counts in SWAT" columns indicate the number of inter-domain steps taken between the given domains in the dataset. The entire process is equivalent to including ("# Given Domains - 1") × ("# Inter-domain counts in SWAT + 1") + 1 training step, which includes self-training of GAN and the encoder $f$ and classifier $g$. For example, with four domains and three intermediate steps, the total number of training steps is calculated as (4 - 1) × (3 + 1) + 1 = 13 small steps.
@@
[…]
6 $\mathbf{99.6\pm 0.0}$ $9.4\pm 0.3 99$$2\pm 0.5 99.$$4\pm 0.1 99.$$9.5\pm 0.1$
@@
It is important to highlight that the highest accuracy points are typically found in the upper-right and lower-left corners of the table. This suggests that as the number of given domains increases, the SWAT tends to become more complete, eliminating the need for additional intermediate steps to refine the flow. This observation demonstrates that our method of constructing flows matching between domains is particularly effective when only a few domains are given, and the sliding window adversarial training is highly effective all the time.
Appendix B Ablation Study on Least Confidence
@@
Appendix C Computational Cost Comparison
Table 9: Running time and peak GPU memory usage (on an RTX 4090) for SWAT, GOAT, and GST (0 inter-domain step) across four benchmarks with 4 given domains and 2 inter-domain step counts.
Method Rotated MNIST Color MNIST Portraits CoverType
SWAT 4 min 56 s / 4488 MB 4 min 59 s / 4488 MB 13 s / 5636 MB 4 min 22 s / 660 MB
GOAT 5 min 38 s / 1886 MB 5 min 43 s / 1392 MB 20 s / 1632 MB 1 min 51 s / 586 MB
GST (0-step) 59 s / 1884 MB 1 min 1 s / 1884 MB 4 s / 990 MB 31 s / 612 MB
Table 9 compares the running time and peak GPU memory consumption of SWAT against GOAT and GST (with zero inter-domain adaptation steps) on an RTX 4090. SWAT incurs only a modest overhead, due to its additional generators and discriminators, while delivering superior task performance. Notably, SWAT matches or outperforms GOAT in both speed and memory usage on most datasets (e.g., 4 min 56 s/4488 MB vs. 5 min 38 s/1886 MB on Rotated MNIST), demonstrating its practical feasibility for large-scale domain adaptation.
Appendix D Limitations
While the proposed Sliding Window Adversarial Training (SWAT) framework effectively addresses challenges in gradual domain adaptation, it has several limitations. It assumes smooth and gradual domain shifts, which may not hold in cases of abrupt or discontinuous changes, potentially impairing adaptation performance. Additionally, SWAT requires access to intermediate unlabeled domains along the domain shift trajectory, which may be limited or unavailable in practical scenarios. The computational cost also grows with the number of intermediate domains and Inter-domain counts, posing challenges for large-scale or high-dimensional data despite memory-efficient design. Although ablation studies highlight sensitivity to hyperparameters and architectures, further evaluation is needed to assess robustness under noisy, biased, or diverse real-world conditions. Future work should aim to relax domain shift assumptions and improve computational efficiency to enhance the applicability of SWAT in broader settings.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the motivation (issues with steep domain shifts in UDA), the problem setup (Gradual Domain Adaptation), and the paper’s proposed solution (Sliding Window Adversarial Training). The stated contributions—including continuous adaptation via adversarial flow, sliding-window alignment, and superior performance on GDA benchmarks—are consistently supported by both theoretical formulations and empirical results in the main body of the paper. See Section 1 (Introduction) for evidence.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper includes a dedicated Limitations section in Appendix D that explicitly discusses key constraints of the work. It highlights known issues with existing Gradual Domain Adaptation (GDA) methods, such as error accumulation from discrete domain partitions and mode collapse due to linear interpolation assumptions, which SWAT is designed to address. Nonetheless, the paper acknowledges that SWAT’s performance depends on assumptions like smooth domain transitions (Section 3) and recognizes practical trade-offs in computational cost and efficiency (Appendix C).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper clearly states all assumptions underlying the theoretical results in Section 3 and 4. All theorems, lemmas, and formulas are properly numbered and referenced. Formal proofs are provided in the supplemental material, accompanied by proof sketches and intuition in the main paper. The proofs reference all necessary intermediate results and are complete and rigorous.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Focusing on the methodology, the architecture and implementation details are carefully elaborated. The paper provides clear descriptions of the model components (generators, discriminators, classifiers), training objectives (e.g., adversarial, cycle-consistent, and self-training losses), and the sliding window mechanism (Section 4). Additionally, Appendix A.1 outlines the network architectures, training configurations, and hardware used. These elements make the experimental results reproducible even without direct access to the code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The paper utilizes publicly available datasets and provides detailed experimental instructions. Most of the code necessary to reproduce the main experiments is included in the supplementary materials. Furthermore, the complete code for the key methods will be released publicly after acceptance to promote transparency and support future research.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper clearly specifies the experimental settings, including dataset splits, intermediate domain construction, and training protocols (Section 5.1). Hyperparameters such as learning rates, optimizer type (Adam), batch sizes, and number of inter-domain steps are detailed in Appendix A.1. The paper also explains confidence-based sample filtering and provides information about hardware used for training. These details enable a full understanding of the experimental results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper reports error bars and confidence intervals for the main experimental results, such as accuracy means ± standard deviations across 5 independent runs (see Tables 2, 5–8). The variability captured includes randomness in training initialization and dataset sampling. The paper explicitly states the use of 68% confidence intervals (approximately ±1 standard deviation). This statistical reporting supports the reliability of the results and the conclusions drawn.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper reports the compute resources used for experiments, including NVIDIA RTX 4090 GPUs, along with running time and peak GPU memory usage for different datasets and methods (Appendix C, Table 9). This information covers the computational cost of training and evaluation for SWAT and baseline methods. While preliminary or failed experiments are not explicitly discussed, the disclosed resources are sufficient to reproduce the reported results.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conforms fully to the NeurIPS Code of Ethics. The paper does not involve sensitive personal data or human subjects, ensures fairness by evaluating across multiple datasets and domain shifts, and does not promote harmful or discriminatory practices. Anonymity is preserved throughout the submission. Ethical considerations related to domain adaptation and machine learning practices have been duly followed.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper discusses both potential positive and negative societal impacts in the conclusion (Section 6). It highlights that SWAT can enhance the robustness and reliability of machine learning systems in real-world domains such as healthcare, autonomous driving, and environmental monitoring. It also acknowledges potential risks, including the reinforcement of biases and privacy concerns if domain adaptation methods are misused or deployed without safeguards.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not involve the release of high-risk models (e.g., large pretrained language models or generative models) or scraped datasets. It primarily focuses on domain adaptation using publicly available datasets and standard architectures, which pose minimal risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The paper uses publicly available datasets such as Rotated MNIST, CIFAR-10C, CIFAR-100C, Portraits, and Cover Type, all of which are properly cited in the references. These datasets are used in accordance with their respective licenses. Additionally, baselines and related methods are cited with appropriate attribution, and no proprietary or scraped data assets are used in this work.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The paper introduces a new GitHub repository that implements the proposed SWAT method, along with necessary scripts and documentation for reproducing the main experimental results. The repository builds upon existing libraries such as autoattack and robustbench, which are properly cited and used in accordance with their licenses. The new asset is well documented and will be made publicly available upon acceptance, with usage instructions provided in the supplementary material.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve any crowdsourcing or research with human subjects. All experiments are conducted using publicly available datasets and automated machine learning models.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve any research with human subjects or crowdsourcing activities. All experiments were conducted using publicly available datasets and do not require IRB approval.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The research does not involve the use of large language models (LLMs) as an important, original, or non-standard component of the core methodology. Any language model use was limited to general writing assistance and did not influence the scientific content, experiments, or technical contributions of the paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.